# mode = titleGraph Filtering for Multimodal Recommendation

### v2

@@ 5.4 Ablation Study (RQ3)
  • $M$-GF-f : setting of tunable hyperparameter $a_{k}$ in Eq. (11) and rather uses the fixed coefficients as in Park et al. (2024, 2025); Kim et al. (2025).
@@
    As evidenced by the results of both MM-GF-t and MM-GF-tv, the gain of further leveraging the visual modality over MM-GF-t is indeed marginal; such a pattern was also comprehensively discussed in the earlier study in Zhou et al. (2023a).
@@
    The performance degradation in MM-GF-pf demonstrates that our modality-specific polynomial filter design contributes to the improvement of performance. Moreover, handling the bound of frequencies is essential for accurate predictions, as MM-GF-f leads to a considerable drop in performance.
Table 6: Performance comparison among MM-GF and competitors in cold-start setting. The best and second-best performers are highlighted in bold and underline, respectively.
@@
Method Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@10 NDCG@10 Recall@20 NDCG@20
@@ 5.5 Analysis in Cold-Start Settings (RQ4)
While high sparsity in graphs leads to technical challenges for model training, leveraging additional features in MRSs can enhance the recommendation accuracy in sparse conditions, as demonstrated by cold-start experiments Zhou et al. (2023a). In this study, we regard the cold-start users as those who have rated equal to or fewer than 5 items. Hence, we only use such cold-start users on three benchmark datasets for model inference. The performance comparison among MM-GF and seven state-of-art multimodal recommendation competitors in cold-start settings is summarized in Table 6.555Since the datasets in cold-start settings differ from the original ones, we determine the optimal hyperparameters for each dataset under these settings. Our observations are made as follows:
@@
00.50.5111.51.5220.010.010.020.020.030.030.040.040.050.05(a) Varying $\beta$ and $\gamma$ on Baby. $\beta$$\gamma$
00.50.5111.51.5220.050.050.0550.0550.060.06(b) Varying $\beta$ and $\gamma$ on Sports. $\beta$$\gamma$
00.50.5111.51.5220.030.030.040.040.050.05(c) Varying $\beta$ and $\gamma$ on Clothing. $\beta$$\gamma$
2-21-1011220.010.010.020.020.030.030.040.040.050.05(d) Varying $a_{k}$ on Baby. $a_{1}$$a_{2}$$a_{3}$
2-21-1011220.020.020.030.030.040.040.050.050.060.06(e) Varying $a_{k}$ on Sports. $a_{1}$$a_{2}$$a_{3}$
2-21-1011220.020.020.030.030.040.040.050.05(f) Varying $a_{k}$ on Clothing. $a_{1}$$a_{2}$$a_{3}$
@@
5.6 Robustness to Noisy Multimodal Features (RQ5)
We analyze the sensitivity of noise that often occurs in real-world scenarios Zhou et al. (2023a). In MRSs, multimodal features are vulnerable to noise due to various factors such as embedding inaccuracies or inconsistencies in data collection processes. To validate the robustness of MM-GF to such noisy multimodal features, we characterize the noise $n$ as $\tilde{X}_{m}^{n}=\tilde{X}_{m}+n$, where $n$ follows $\mathcal{N}(0,\sigma_{m}^{2})$ for the standard deviation $\sigma_{m}$. We define six different levels of noise according to different levels of the standard deviation: level 0 corresponds to the case where there is no noise, which represents the original datasets; level 1 corresponds to the noise equivalent to $10\%$ of the standard deviation of each feature embedding; and as noise gets gradually added, level 5 corresponds to the noise that is twice the standard deviation of each feature embedding.
Figure 5.6 clearly shows that as the multimodal information incorporates more noise, all the competitors exhibit a decreasing trend in NDCG@$20$. In particular, MGCN Yu et al. (2023) purifies the modality features to prevent noise contamination, which is a method directly designed for noise removal. In contrast, MM-GF consistently reveals the best performance in MRSs through only matrix computations without any noise-purification design. In other words, MM-GF exhibits robustness to noise with a simple yet effective multimodal feature process.
5.7 Hyperparameter Sensitivity
@@
    As shown in Figure 7a, there is a monotonically decreasing pattern with increasing $\gamma$. That is, using visual features in MM-GF is rather harmful for multimodal recommendations, which is consistent with the earlier work in Zhou et al. (2023a) that confirmed that not all multimodal information contributes to performance improvement.
@@
    While the optimal values of hyperparameters differ across datasets and necessitate a broad search range, MM-GF is inherently less sensitive to hyperparameter choices and allows for rapid tuning. This is in contrast to training-based methods Wei et al. (2020); Zhang et al. (2021); Yu et al. (2023); Zhou and Shen (2023), which often demand costly and time-consuming optimization procedures. Our model’s simplicity enables efficient deployment even with minimal tuning effort.
@@ 6 Conclusions and Outlook
In this paper, we proposed MM-GF, the first training-free multimodal recommendation method grounded in the concept of GF, aiming for both efficiency and accuracy. MM-GF effectively adjusted the spectral bounds of frequencies, alongside a flexible and spectrally-aware graph filtering strategy. By integrating modality-specific similarity graphs with a tunable polynomial graph filter, our MM-GF method was shown to enable precise frequency control and robust fusion of heterogeneous information. Extensive experimental evaluations demonstrated not only the remarkably fast runtime of MM-GF but also the superior recommendation accuracy of MM-GF in diverse challenging scenarios, including cold-start conditions and resilience to noisy features. Avenues of future research include the design of scalable GF methods that accommodate large-scale multi-modal feature data.
@@ Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by Korea government (MSIT) under Grants RS-2021-NR059723 and RS-2023-00220762.

### v3

@@ 5.4 Ablation Study (RQ3)
  • MM-GF-f : rather uses the fixed coefficients $a_{k}$ in Eq. (11) as in DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25;
  • MM-GF-h : constructs the model without spectral bound adjustment in Eq. (11), i.e., directly applies the polynomial filter to the original spectrum as in DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25.
@@
    As evidenced by the results of both MM-GF-t and MM-GF-tv, the gain of further leveraging the visual modality over MM-GF-t is indeed marginal; such a pattern was also comprehensively discussed in the earlier study in zhou2023comprehensive.
@@
    The performance degradation in MM-GF-pf demonstrates that our modality-specific polynomial filter design contributes to the improvement of performance. Moreover, handling the bound of frequencies is essential for accurate predictions, as MM-GF-f and MM-GF-h lead to a considerable drop in performance.
Table 6: Performance comparison among MM-GF and competitors in cold-start setting. The best and second-best performers are highlighted in bold and underline, respectively. Here, R@$K$ and N@$K$ indicates Recall@$K$ and NDCG@$K$, respectively.
@@
Method R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20
@@ 5.5 Analysis in Cold-Start Settings (RQ4)
While high sparsity in graphs leads to technical challenges for model training, leveraging additional features in MRSs can enhance the recommendation accuracy in sparse conditions, as demonstrated by cold-start experiments zhou2023comprehensive. In this study, we regard the cold-start users as those who have rated equal to or fewer than 5 items. Hence, we only use such cold-start users on three benchmark datasets for model inference. The performance comparison among MM-GF and seven state-of-art multimodal recommendation competitors in cold-start settings is summarized in Table 6.555Since the datasets in cold-start settings differ from the original ones, we determine the optimal hyperparameters for each dataset under these settings. Our observations are made as follows:
@@
  3. (iii)
    From a modeling perspective, this robustness can be attributed to the graph-filtering nature of MM-GF. Since trainable GNN-based recommenders DBLP:conf/aaai/HeM16; DBLP:conf/mm/WeiWN0C20; DBLP:conf/mm/Zhang00WWW21; DBLP:conf/www/ZhouZLZMWYJ23; DBLP:conf/mm/ZhouS23; DBLP:conf/mm/Yu0LB23 rely on parameter updates that are directly driven by historical interactions, items with very few edges receive weak or unstable updates and tend to be poorly represented in cold-start regimes. In contrast, MM-GF constructs modality-specific item–item similarity graphs and then applies a polynomial graph filter to these graphs; even for cold-start items with sparse interactions, their representations are smoothed toward those of spectrally similar and well-connected neighbors. Consequently, the effective information available to such items is enriched through GF, which explains why MM-GF maintains relatively stable performance under sparse (cold-start) conditions.
5.6 Robustness to Noisy Multimodal Features (RQ5)
We analyze the sensitivity of noise that often occurs in real-world scenarios zhou2023comprehensive. In MRSs, multimodal features are vulnerable to noise due to various factors such as embedding inaccuracies or inconsistencies in data collection processes. To validate the robustness of MM-GF to such noisy multimodal features, we characterize the noise $n$ as $\tilde{X}_{m}^{n}=\tilde{X}_{m}+n$, where $n$ follows $\mathcal{N}(0,\sigma_{m}^{2})$ for the standard deviation $\sigma_{m}$. We define six different levels of noise according to different levels of the standard deviation: level 0 corresponds to the case where there is no noise, which represents the original datasets; level 1 corresponds to the noise equivalent to $10\%$ of the standard deviation of each feature embedding; and as noise gets gradually added, level 5 corresponds to the noise that is twice the standard deviation of each feature embedding.
Figure 6 clearly shows that as the multimodal information incorporates more noise, all the competitors exhibit a decreasing trend in NDCG@$20$. In particular, MGCN DBLP:conf/mm/Yu0LB23 purifies the modality features to prevent noise contamination, which is a method directly designed for noise removal. In contrast, MM-GF consistently reveals the best performance in MRSs through only matrix computations without any noise-purification design. In other words, MM-GF exhibits robustness to noise with a simple yet effective multimodal feature process.
From a spectral viewpoint, this robustness can be explained by the design of MM-GF. Additive Gaussian noise on multimodal feature embeddings mainly manifests as high-frequency components DBLP:journals/spm/ShumanNFOV13; DBLP:conf/icml/YuQ20, whereas the proposed polynomial graph filter behaves as a controlled low-pass operator on the bounded spectrum. By attenuating high-frequency fluctuations while preserving the dominant low-frequency structure, MM-GF effectively suppresses noise in multimodal features without any explicit denoising module. Moreover, since our method is training-free and does not learn a large number of parameters from noisy input, it is less prone to overfitting to noise than heavily parameterized GNN-based models, which further contributes to its stable performance under noisy conditions.
(a) Baby
(b) Sports
(c) Clothing
Figure 7: Frequency response functions tuned on three datasets with respect to the graph frequency $\lambda$.
(a) Unbounded 1st-order polynomial LPF
(b) Bounded 1st-order polynomial LPF
(c) Unbounded 2nd-order polynomial LPF
(d) Bounded 2nd-order polynomial LPF
(e) Unbounded 3rd-order polynomial LPF
(f) Bounded 3rd-order polynomial LPF
Figure 8: Eigenvalue distributions under different polynomial LPFs, with left figures showing unbounded spectra without our spectral regularization and right ones illustrating well-bounded spectra after applying the proposed normalization for 1st-, 2nd-, and 3rd-order polynomial filters, respectively.
5.7 Spectral Analysis (RQ6)
We examine how the polynomial graph filter of MM-GF operates under different spectral conditions across datasets. Figure 7 illustrates the frequency response functions derived from the user–item interaction graphs of three datasets with respect to the graph frequency $\lambda$. Owing to their distinct eigenvalue distributions, each dataset induces a different spectral range $0,\lambda^{*}]$, which leads the filter to adopt dataset-specific attenuation patterns. This demonstrates that the proposed polynomial filter adapts its spectral emphasis to the underlying graph structure rather than imposing a fixed response, thereby highlighting that the adaptable polynomial coefficients of MM-GF play a central role in shaping its behavior under different spectral conditions.
Additionally, we analyze the eigenvalue convergence induced by the proposed spectral regularization. Figure 8 compares the eigenvalue distributions before and after applying our spectral normalization in Eq. (11). When the polynomial filters are applied without our regularization (see Figures 8(a), 8(c) and 8(e)), the resulting spectra easily escape the desirable range $0,\lambda^{\ast}]$. This issue becomes dramatically evident for higher-order filters. For example, the 3rd-order polynomial LPF $h(\lambda)=1-\lambda^{3}$ produces eigenvalues approaching $1,000$, which makes the filter unstable and undermines any meaningful interpretation of the spectral response. Such divergence clearly indicates that unbounded polynomial filters cannot operate reliably on real-world user–item graphs, and the spectral domain quickly becomes distorted.
In contrast, Figures 8(b), 8(d), and 8(f) show the spectra obtained after applying our spectral regularization, which rescales the graph operator so that all polynomial filters (1st-, 2nd-, and 3rd-order polynomials) produce eigenvalues that remain strictly within a compact interval. This boundedness demonstrates the convergence stability of the proposed method; that is, regardless of polynomial orders, the filtered operator does not amplify spectral components beyond the specific range. Consequently, the proposed polynomial graph filters preserve their intended low-pass characteristics and remain interpretable in the spectral domain.
@@ 5.7 Spectral Analysis (RQ6)
5.8 Hyperparameter Sensitivity (RQ7)
@@
    As shown in Figure 9a, increasing the weight $\gamma$ of the visual modality leads to a consistent performance drop, suggesting that visual features may introduce noise rather than benefit recommendation accuracy. This observation aligns with prior findings zhou2023comprehensive, which report that incorporating all modalities does not universally improve performance and that the usefulness of visual features varies depending on dataset characteristics and feature quality.
@@
    While the optimal values of hyperparameters differ across datasets and necessitate a broad search range, MM-GF is inherently less sensitive to hyperparameter choices and allows for rapid tuning. This is in contrast to training-based methods DBLP:conf/mm/WeiWN0C20; DBLP:conf/mm/Zhang00WWW21; DBLP:conf/mm/Yu0LB23; DBLP:conf/mm/ZhouS23, which often demand costly and time-consuming optimization procedures. Our model’s simplicity enables efficient deployment even with minimal tuning effort.
Table 7: The performance according to the polynomial order $K$ for the user–item interaction graph with respect to NDCG@20. The best performer is highlighted in bold.
Dataset $K=1$ $K=2$ $K=3$
Baby 0.0465 0.0499 0.0522
Sports 0.0547 0.0581 0.0593
Clothing 0.0448 0.0467 0.0513
Furthermore, to validate the choice of the polynomial order, we conduct an additional experiment by varying the orders for the similarity graph constructed from user–item interactions. In case of multimodal information, we set the polynomial order to 1, as our experiments indicate that further increasing the orders not only introduces additional hyperparameters but also leads to performance degradation. We vary the polynomial order as $K\in\{1,2,3\}$ while keeping all other settings fixed. As summarized in Table 7, increasing $K$ from 1 to 3 yields consistent improvements in NDCG@20 across all three datasets. We do not extend the analysis to $K>3$, since higher order introduce additional polynomial coefficients and require more matrix multiplications, resulting in substantially higher computational cost and, in practice, undermine the efficiency benefits expected from a training-free framework.
@@ 6 Conclusions and Outlook
In this paper, we proposed MM-GF, the first training-free multimodal recommendation method grounded in the concept of GF, aiming for both efficiency and accuracy. MM-GF effectively adjusted the spectral bounds of frequencies, alongside a flexible and spectrally-aware GF strategy. By integrating modality-specific similarity graphs with a tunable polynomial graph filter, our MM-GF method was shown to enable precise frequency control and robust fusion of heterogeneous information. Extensive experimental evaluations demonstrated not only the remarkably fast runtime of MM-GF but also the superior recommendation accuracy of MM-GF in diverse challenging scenarios, including cold-start conditions and resilience to noisy features. Potential avenues of future research include adaptive mechanisms to make the filtering and modality fusion more data-dependent. Future research in this area in the inclusion of an extension to streaming recommendation, where the training-free nature of MM-GF allows frequent refresh of recommendations under rapidly changing interactions and contents.
Nevertheless, MM-GF has several challenges and limitations. First, since we directly use pre-trained textual and visual embeddings without additional modality-specific pre- or post-processing, the contribution of multimodal information can be limited to some extent when these features are noisy or weak. Second, the static design of MM-GF—where filter coefficients and modality weights are fixed after validation—ensures efficiency and simplicity, but requires manual retuning under shifts in data distribution or modality balance. We leave addressing these issues—for example, by incorporating lightweight adaptive mechanisms or more sophisticated multimodal processing while preserving the training-free spirit—as important directions for future work.
@@ Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2021-NR059723, No. RS-2023-00220762) and by SMEs Technology Innovation Development Program through the Technology Innovation and Promotion Agency (TIPA), funded by Ministry of SMEs and Startups (RS-2024-00511332).